OpenStudy (anonymous):

Find the domain of the following function: f (x) = (3x+5) / (4x^2+9) I know that the denominator cannot be equal to zero, but I'm getting stuck after that

OpenStudy (anonymous):

Sorry determine what value for x would be in the equation 4x^2+9 =0

OpenStudy (anonymous):

X would equal the sqr root of -9/4

OpenStudy (anonymous):

So the domain would be x is not equal to + or - sqr root (9/4)?

OpenStudy (anonymous):

X^2 = -9/4 or as you show x = square root of -9/4. Are either of these possible?

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

No, they will both give an answer of zero

OpenStudy (anonymous):

denominator is never 0 for any real value of x.

OpenStudy (anonymous):

There is no value that will result in a denominator of zero then, correct?

OpenStudy (anonymous):

so domain is all real values.

OpenStudy (anonymous):

Thus x can be all real numbers

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

Neither one of those are possible.

OpenStudy (anonymous):

Yes, @surjithayer . I know that the denominator cannot equal zero. I'm having trouble with finding the domain

OpenStudy (anonymous):

domain is the value of x which can satisfy the given relation. here x can have all real values so domain is all real values .